# Adaptive Modality Balanced Online Knowledge Distillation for Brain-Eye-Computer based Dim Object Detection

### v2

@@ V-C2 Visualization Analysis
The latent representations of the models in the ablation analysis are visualized using the t-distributed stochastic neighbor embedding (t-SNE) approach to further verify the feature extraction and fusion capability of our proposed approach [68]. As depicted in Fig. 10, the feature distributions of all subjects are displayed by different colored dots, with red dots representing the features in the target domain, and dots in other colors representing the features in the non-target class.
As can be seen from Fig. 10, unimodality models such as EfficientNet and MCGRAM exhibit distinct visual and EEG modality characteristics in their feature representations. The visual-based EfficientNet model exhibits a more discrete feature representation, with a large intra-class distance, indicating its superior feature extraction performance but poor generalization performance of the classifier. Conversely, the feature representation of the MCGRAM model based on EEG is more clustered, with a small intra-class distance, indicating weaker feature extraction ability but stronger classifier generalization. As a multimodal fusion model, AMM effectively combines the benefits of vision and EEG modalities, thereby further reducing the inter-class distance while maintaining the feature representation capability. By integrating single-teacher guidance mechanisms into the multimodal fusion model, V-KD and E-KD models exhibit enhanced classification performance and generalization, as reflected in larger inter-class and smaller intra-class differences in feature representation. In addition, by introducing a multi-teacher mechanism, OKD approach, and multimodal AMB approach, the discrimination of feature representation extracted by MKD, MMOKD, and AMBOKD models become more pronounced and the feature representation distances within the same class become closer. This indicates that these mechanisms enhance the capability of multimodal fusion and effectively improve the classifier’s accuracy and generalization.
In conclusion, the effectiveness of the AMBOKD framework and its modules is effectively verified by the ablation study. The proposed AMBOKD approach outperforms other methods in the multimodal fusion task, exhibiting robust discrimination between positive and negative samples, an increase in inter-class distance, a reduction in intra-class distance, and statistically significant experimental findings ($p<0.05$).
@@ V-D Adaptive Modality Balancing Analysis
Figure 12: Sensitivity analysis on AMB module.
Figure 13: The performances of the uni-modality and the unimodality trained with AMBOKD in generalization analysis.
We then analysis the effectiveness of the AMB module and the influence of hyperparameter $\gamma$ on the AMB module.
V-D1 Effectiveness Analysis
We use the dynamic learning progress ratio $R^{DG}_{(t)}$ to assess the effectiveness of the modality equilibrium. The first training data from a 10-fold cross-validation technique with a seed value of 1 was utilized.
As illustrated in Fig. 11, we draw the changes of $R^{DG}_{(t)}$ for each modality within MMOKD, MMOKD-DG, and AMBOKD. In the standard MMOKD approach, despite the intermodal learning capabilities, optimization is impeded by the inherent characteristics of each modality. Notably, the EEG modality progresses at a slower rate compared to the Visual and Fusion modalities, which exhibit quicker optimization speeds. This discrepancy in training speeds results in a suboptimal mutual learning effect among the modalities. The Dynamic Gradient (DG) block is able to addresses this issue by dynamically adjusting the training gradients according to each modality’s learning progression. With this adjustment, MMOKD-DG improves the optimization speed and the final optimization degree of each modality, especially EEG modality. AMBOKD further employs a dynamic weight modulation module tailored to KL loss, optimizing the faster-learning Visual and Fusion modalities initially, while also supporting sustained development in the EEG modality. This strategy ensures that all three modalities are optimally and uniformly enhanced throughout the training process, achieving superior overall performance.
V-D2 Sensitivity Analysis
The sensitivity control of the AMB module is based on the hyperparameter $\gamma$, which can be found in Eq. (14). As shown in Fig. 12, the AMBOKD method with $\gamma=3$ achieves the highest AUC in session 1 and second-highest AUC in session 2, resulting in an improved classification performance for imbalanced samples.
V-E Generalization Analysis
To further analysis the generalization performance of the AMBOKD method, we conduct experiments on the algorithm level and the data level.
V-E1 Generalization Analysis on Algorithm Level
Comparative experiments are conducted by replacing the unimodal backbone networks used in the EEG and visual encoders. With these experiments, we are able to further validate the generalization capability of the proposed AMB module in multimodal learning but also its ability on improving unimodality performance. We test the following combinations: EEGNet + EfficientNet, TSception + EfficientNet, MCGRAM + MobileNet, MCGRAM + ResNet50, EEGNet + MobileNet, and TSception + ResNet50.
TABLE V: The subject-independent reuslts of the baseline and state-of-art methods in generalization analysis
@@ V-E1 Generalization Analysis on Algorithm Level
As shown in Fig. 13, the unimodality models trained with AMBOKD approach consistently achieve higher performances than the models trained by their own in all combinations. Some of the uni-modal models show significant performance improvements over their original counterparts after training under AMBOKD. For example, EEGNet trained under AMBOKD outperforms the original results in AUC by 7.7 %, ResNet50 trained under AMBOKD shows an improvement in AUC of 9.3 % in AUC (see Fig. 13 (a)$\sim$(c)). These findings indicate the superiority and generalization capabilities of the AMBOKD in multimodal learning tasks.
V-E2 Generalization Analysis on Data Level
We verify the generalization of the AMBOKD method by conducting transfer experiments under few-sample conditions. In this experiment, we utilize the optimally saved model that is specifically trained for armored vehicle recognition as detailed in Section  V-B1. This model is further refined using 30 real-world scene samples and is deployed to detect vehicle targets within these real scenes. The comparative results, presented in Table V, demonstrate that our methods achieved superior performance in both session 1 and session 2. This case effectively verifies the effectiveness and generalization ability of our method.
@@ V-F Physical Verification
To further verify the pratical and effectiveness of our build system and proposed method, we carry our physical experiments combined with the application scenario of UAV ground control station operators, as illustrated in Fig. 14. During the system experiment, subjects are positoned in front of a computer screen with proper body alignment, focusing their gaze directly on the frontal display while performing the ESSVP paradigm. They are tasked with controlling the system and issuing appropriate commands to the UAV upon detecting abnormal displays on the UAV monitoring interface located to the participant’s left. Concurrently, the ESSVP paradigm temporarily halts the playback of visual stimuli to accommodate this interaction.
@@
In the system verification experiment, the models refined in SectionV-E are employed directly to process the multimodal data and detect the target. As shown in Fig. 15, the AMBOKD method achieves the highest performances in all metrics compare with the baseline and state-of-art methods. This experimental findings demonstrate the effectiveness of the brain-eye-computer based object detection system in real-world settings, indicating robustness and high accuracy in practical tasks.
Figure 15: The results on physical verification experiment.
@@
Further more, we compared our brain-eye-computer based system with traditional manual and computer vision approaches, as shown in Table VI. In the human-based experiment, five subjects were asked to detect and annotate vehicle targets in 60 aerial images. The results indicate variability in detection time and accuracy among subjects, underscoring the inefficiency and unreliability of purely manual methods in complex scenarios. In the computer-based experiment, we adopt the same training setup as the brain-eye-computer based system, a Faster RCNN model with a ResNet50 backbone trained on the public dataset was fine-tuned with 30 samples and then tested on 60 images. Although this method achieve rapid detection capabilities, its accuracy was notably compromised due to the limited training samples. In contrast, our integrated brain-eye-computer system achieved superior performance, registering a total processing time of 245.37 s (240 s for the ESSVP paradigm and 5.37 s for processing). This result not only demonstrates the efficiency of the system, but also its robust generalization capabilities across different test conditions.
VI Conclusion
This study builds a brain-eye-computer based system for object detection in aerial images under few-shot conditions. This system detects suspicious targets in aerial images using a region proposal network. After obtaining the images with the region proposals, it elicits subjects’ ERP signals during target search with the ESSVP paradigm and constructs EEG-image data pairs incorporating eye movement data. These pairs are then recognized using the proposed AMBOKD method. AMBOKD fully extracts and fuses crucial information from EEG and visual modality features, facilitates end-to-end mutual learning, and improves adaptive multimodal interaction capability through the AMB module. Experimental results demonstrates the effectiveness and superiority of our method, and its ability in improving the performance of uni-modality models. Lastly, the feasibility and transferability of the AMBOKD method and the brain-eye-computer based system are verified through experiments with practical scenario images. This study opens new possibilities for robust and efficient dim object detection in aerial applications.
@@ VI Conclusion
Human 78.28/18.20 67.33/20.26 72.00/18.71 1026.00/362.00
Computer 23.44/5.49 30.12/6.89 26.36/6.10 3.00/0
Brain-eye-computer 89.26/6.76 90.32/4.01 87.98/6.37 245.37/1.38
In the future, we will further develop real-time online systems across a wider range of scenarios. We aim to design more efficient experimental paradigms and methods that take full advantage of multimodal characteristics, enhancing the effectiveness and adaptability of multimodal fusion methods.

### v3

@@ V-C2 Visualization Analysis
The latent representations of the models in the ablation analysis are visualized using the t-distributed stochastic neighbor embedding (t-SNE) approach to further verify the feature extraction and fusion capability of our proposed approach [55]. As depicted in Fig. 8, the feature distributions of all subjects are displayed by different colored dots, with red dots representing the features in the true target domain, and dots in other colors representing the features in the true non-target domain.
As can be seen from Fig. 8, unimodality models such as EfficientNet and MCGRAM exhibit distinct visual and EEG modality characteristics in their feature representations. The visual-based EfficientNet model exhibits a more discrete feature representation, with a large intra-class distance, indicating its superior feature extraction performance but poor generalization performance of the classifier. Conversely, the feature representation of the MCGRAM model based on EEG is more clustered, with a small intra-class distance, indicating weaker feature extraction ability but stronger classifier generalization. As a multimodal fusion model, AMM effectively combines the benefits of vision and EEG modalities, thereby further reducing the inter-class distance while maintaining the feature representation capability. By integrating single-teacher guidance mechanisms into the multimodal fusion model, V-KD and E-KD models exhibit enhanced classification performance and generalization, as reflected in larger inter-class and smaller intra-class differences in feature representation. Furthermore, by introducing a multi-teacher mechanism, OKD, and the AMB module, the feature representations of MKD, MMOKD, and AMBOKD models exhibit enhanced discrimination and reduced intra-class distances. This indicates that these mechanisms enhance the capability of multimodal fusion and effectively improve the classifier’s accuracy and generalization.
In conclusion, the ablation study confirms the effectiveness of the AMBOKD framework and its modules. The proposed AMBOKD approach outperforms other methods in the multimodal fusion task, exhibiting robust discrimination between positive and negative samples, an increase in inter-class distance, a reduction in intra-class distance, and statistically significant experimental findings ($p<0.05$).
@@ V-D Adaptive Modality Balancing Analysis
In this section, we use the dynamic learning progress ratio $R^{DG}_{(t)}$(Eq. 13), which is capable of showing the dynamic optimization level of each modality to assess the effectiveness of the AMB module.
As illustrated in Fig. 9, we draw the changes of $R^{DG}_{(t)}$ for each modality within MMOKD, MMOKD-DG, and AMBOKD during the training process on the ESSVP dataset, which can effectively indicate the real-time optimization degree of each modality. Although the standard MMOKD approach supports intermodal learning, its optimization is hindered by the inherent characteristics of each modality. This discrepancy in optimization results in a suboptimal mutual learning effect among the modalities, convergent feature vectors cannot be learned between modalities. The Dynamic Gradient (DG) block is able to address this issue by dynamically adjusting the training gradients according to each modality’s learning progression. With this adjustment, MMOKD-DG improves the optimization speed and the final optimization degree of each modality. AMBOKD further employs a dynamic weight modulation module tailored to KL loss, guaranteeing the further optimization of higher performing modalities, while also supporting sustained development in the slower modality. This strategy ensures that all three modalities are optimally and uniformly enhanced throughout the training process, learning more consistent features and achieving superior overall performance.
TABLE V: The subject-independent results of the baseline and state-of-art methods in generalization analysis
@@
V-E Generalization Analysis
To further analyze the generalization performance of the AMBOKD method in different scenarios, we conduct transfer experiments from simple scenes to complex scenes under few-sample conditions. (the scene images are shown in Fig. 3).
In this experiment, we utilize the optimally saved model trained in the subject-independent experiment. This model is further refined using 30 real-world scene samples and is deployed to detect vehicle targets within these real scenes. The comparative results, presented in Table V, demonstrate that the accuracy of our method decreases in complex scenes after transfer experiments, but it is still higher than all other comparison methods in both session 1 and session 2. These results confirm the effectiveness and generalization ability of our method.
Figure 11: The results on physical verification experiment.
@@ V-F Physical Verification
To verify the practical effectiveness of our built system and the proposed method, we carry out our physical experiments combined with the application scenario of UAV ground control station operators, as illustrated in Fig. 10. During the system experiment, subjects are positioned in front of a computer screen with proper body alignment, focusing their gaze directly on the frontal display while performing the ESSVP paradigm. They are tasked with controlling the system and issuing appropriate commands to the UAV upon detecting abnormal displays on the UAV monitoring interface located to the participant’s left. Concurrently, the ESSVP paradigm temporarily halts the playback of visual stimuli to accommodate this interaction.
@@
In the system verification experiment, the models refined using 30 real-world scene samples are employed directly to process the multimodal data and detect the target. As shown in Fig. 11, the AMBOKD method achieves the highest performances in all metrics compared with the baseline and state-of-the-art methods. These experimental findings demonstrate the effectiveness of the brain-eye-computer based object detection system in real-world settings, indicating robustness and high accuracy in practical tasks.
@@
We further compared our brain-eye-computer based system with traditional manual and computer vision approaches, as shown in Table VI. In the human-based experiment, five subjects with no prior expertise in aerial image analysis were recruited to detect and annotate vehicle targets in 60 aerial images, ensuring a fair baseline for manual detection. Each subject underwent a 10-minute training session to familiarize themselves with the annotation guidelines, and the experiment was conducted in a controlled environment with consistent lighting and display settings to ensure uniformity. The results indicate significant variability in detection time and accuracy among subjects, with a mean accuracy of 78.20±18.20% and a detection time of 1026.00±360.00s (as shown in Table VI). This high variance highlights the inconsistency of human performance in such tasks, underscoring the inefficiency and unreliability of purely manual methods in complex scenarios, as stable performance is challenging without automated assistance. In the computer-based experiment, we adopted the same training setup as the brain-eye-computer based system: a Faster RCNN model with a ResNet50 backbone, trained on the public dataset, was fine-tuned with 30 samples and then tested on 60 images. Although this method achieves rapid detection capabilities, its accuracy was notably compromised due to the limited training samples. In contrast, our integrated brain-eye-computer system achieved superior performance, registering a total processing time of 245.37s (240s for the ESSVP paradigm and 5.37s for data analysis). This result not only demonstrates the efficiency of the system but also its robust generalization capabilities across different test conditions.
@@
[…]
VI Conclusion and Future Work
This study builds a brain-eye-computer based system for object detection in aerial images under few-shot conditions. This system uses a region proposal network to detect suspicious targets in aerial images. After obtaining the images with the region proposals, it elicits subjects’ ERP signals during target search with the ESSVP paradigm and constructs EEG-image data pairs incorporating eye movement data. These pairs are then recognized using the proposed AMBOKD method. AMBOKD fully extracts and fuses crucial information from EEG and visual modality features, facilitates end-to-end mutual learning, and improves adaptive multimodal interaction capability through the AMB module. Experimental results demonstrate the effectiveness and superiority of our method, as well as its ability to enhance unimodal model performance. Lastly, the feasibility and transferability of the AMBOKD method and the brain-eye-computer based system are verified through experiments with practical scenario images. This study opens new possibilities for robust and efficient dim object detection in aerial applications.
Despite its promising results, the proposed method has several limitations that should be addressed in future research. First, the current system is designed and evaluated in offline settings, which may limit its real-time applicability in fast-paced aerial detection tasks. Future efforts will focus on developing real-time online systems to improve response time and practical utility across a broader range of scenarios. Second, the system’s performance may vary based on the subject’s familiarity and engagement with the ESSVP paradigm, necessitating more adaptive and efficient experimental paradigms to reduce inter-subject variability. Additionally, to further reduce experimental and application costs, future work will explore system experiments based on low-cost EEG devices with fewer channels. Finally, while the study recruited a controlled sample to ensure consistency, the homogeneity of participants may affect generalizability. Expanding the dataset with more diverse subjects and aerial images will help enhance robustness and applicability in real-world scenarios.